# VIPeR: Visual Incremental Place Recognition with Adaptive Mining and Lifelong Learning

### v1

@@ Abstract
Visual place recognition (VPR) is an essential component of many autonomous and augmented/virtual reality systems. It enables the systems to robustly localize themselves in large-scale environments. Existing VPR methods demonstrate attractive performance at the cost of heavy pre-training and limited generalizability. When deployed in unseen environments, these methods exhibit significant performance drops. Targeting this issue, we present VIPeR, a novel approach for visual incremental place recognition with the ability to adapt to new environments while retaining the performance of previous environments. We first introduce an adaptive mining strategy that balances the performance within a single environment and the generalizability across multiple environments. Then, to prevent catastrophic forgetting in lifelong learning, we draw inspiration from human memory systems and design a novel memory bank for our VIPeR. Our memory bank contains a sensory memory, a working memory and a long-term memory, with the first two focusing on the current environment and the last one for all previously visited environments. Additionally, we propose a probabilistic knowledge distillation to explicitly safeguard the previously learned knowledge. We evaluate our proposed VIPeR on three large-scale datasets, namely Oxford Robotcar, Nordland, and TartanAir. For comparison, we first set a baseline performance with naive finetuning. Then, several more recent lifelong learning methods are compared. Our VIPeR achieves better performance in almost all aspects with the biggest improvement of $13.65\%$ in average performance.
@@ I Introduction
Visual place recognition (VPR) aims to recognize a previously visited place by giving visual observations like images. It is a core component in simultaneous localization and mapping (SLAM) systems [1] to ensure the robust operation of autonomous agents and augmented/virtual reality (AR/VR) systems in large-scale environments. When solving the VPR task, most of the existing approaches [2, 3, 4] cast it as a retrieval task. The solution then typically involves a two-step process, with step one generating global descriptors from local features and step two matching the descriptors to the ones in a database of place-tagged descriptors.
Traditionally, VPR methods use hand-craft local features or semantic features and construct global descriptors with vector of locally aggregated descriptors (VLAD) [5], graph-based descriptors [6], etc. More recently, following the success of NetVLAD [2], there have been notable advancements in end-to-end VPR methods in terms of accuracy [3, 4] and robustness [7, 8]. However, such a boost in performance comes at a cost. As the network architecture used in modern deep learning-based VPR methods getting much more complex, for example, KPConv [9] used in AEGIS-Net [4] and Transformer [10] used in $R^{2}$ormer [11], a laborious pre-training process that commonly takes days or even weeks becomes inevitable. In addition, because the trained models have a strong independent and identical distribution (i.i.d.) assumption on the test data, the performance drops considerably when the test data is collected from an unseen environment that doesn’t follow the i.i.d. assumption.
@@
However, when deploying the VPR-equipped autonomous agents and AR/VR systems in real-world scenarios, it is difficult to collect sufficient data that covers all potential environments beforehand to pre-train the deep learning-based models. Such a gap between the advance in deep learning-based VPR methods and the practical needs of autonomous agents and AR/VR systems leads to the fact that recent practical-oriented systems still prefer the traditional VPR methods [12, 13]. In this paper, we aim to close this gap by enabling the lifelong learning ability of the deep learning-based VPR method. In particular, we present VIPeR, a visual incremental place recognition method that can adapt to the newly observed environment during deployment while reserving its knowledge of previously visited environments. Fig. 1 illustrates the core idea of the proposed VIPeR.
@@
The idea of enabling lifelong learning for deep neural networks is not new. Also known as continual learning or incremental learning, it is designed to overcome the notorious catastrophic forgetting [14] when learning from sequential inputs. Despite that lifelong learning has been widely studied [15] and applied to many computer vision [16] or natural language processing [17] tasks, there is only a handful of works attempting to apply it to the place recognition task. AirLoop [18] and InCloud [19], as the pioneers in lifelong place recognition, both tackle the catastrophic forgetting with a simple memory structure and employ knowledge distillation [20] to constrain the training. AirLoop further employs the memory aware synapses (MAS) [21] for regularization.
@@
Although achieving promising performance, there is still plenty of room for improvement. As pointed out in CCL [22], the naive triplet loss used in AirLoop [18] and InCloud [19] inevitably introduces bias when learning the global representation of various places. Such a phenomenon becomes even severe when hard-mining is applied to the training (see Table II). In addition, the naive choice of a simple memory structure substantially hinders the power of the rehearsal strategy. Finally, the deterministic knowledge distillation used in AirLoop and InCloud is very sensitive to the outliers in the data, leading to potential performance degradation.
@@
Targeting these issues, we first introduce an adaptive mining scheme for triplet loss in metric learning. Our adaptive mining scheme can adaptively choose hard mining or soft mining, achieving a balanced performance between a single environment and multiple environments. Focusing on lifelong learning, we then draw inspiration from human memory systems and design a novel memory bank to store memories of multiple environments. Our memory bank is made of three components. For the current environment, we have a sensory memory bank and a working memory bank. For previously visited environments, we have another long-term memory bank. Finally, to make the model more robust against outliers, we extend the deterministic knowledge distillation used in AirLoop [18] and InCloud [19] to a probabilistic one. We conduct extensive experiments on three large-scale datasets and our proposed VIPeR demonstrates superior performance when compared to other state-of-the-art (SoTA) lifelong place recognition methods.
@@
  • We design a novel, brain-inspired memory bank for rehearsal, alleviating the effects of catastrophic forgetting;
@@
  • We conduct extensive experiments on three publicly available datasets and demonstrate SoTA performance.
@@
II-A Lifelong Learning
Lifelong learning targets the sequential learning setting and focuses on improving the adaptability and generalizability of the deep learning models. However, naive solutions like finetuning on the new tasks make the model quickly forget the knowledge about previous tasks, resulting in catastrophic forgetting [14]. Addressing this issue, most existing methods can be categorized into rehearsal-based approaches and regularization-based approaches [16].
The core of the rehearsal-based approach is a memory buffer that either explicitly or implicitly stores samples from previous tasks. Then, when the model is trained on a new task, the samples stored in the buffer join the samples of the current task to refresh the model’s memory about previous tasks. Exemplary works in this approach are iCaRL [23], gradient episodic memory (GEM) [24], and deep generative replay (DGR) [25], which tackle the image classification task with iCaRL and GEM explicitly stores a small number of samples per class for rehearsal and DGR accompanies an additional generative model to enable implicit rehearsal.
@@ II-B Learning-based Place Recognition
NetVLAD [2], as a milestone work in the learning-based VPR method, modified the VLAD [5] descriptor to make it differentiable and enabled end-to-end training of VPR methods. Demonstrating the representation power of the deep learning models, it inspires a plethora of works to promote the performance of deep learning-based VPR methods.
To improve the performance with image inputs, SPE-VLAD [29] explored the spatial pyramid structure of images to make the VLAD descriptor aware of structural information. Addressing the problems of appearance and viewpoint changes, Patch-NetVLAD [30] proposed to combine the advantages of local and global features by fusing multi-scale patch-level features that are computed from NetVLAD residuals. Moving on to 3D inputs like point clouds, PointNetVLAD [31] is a straightforward extension of NetVLAD by replacing the 2D CNN with the PointNet [32] model to handle the input point clouds of fixed sizes. To empower the model to deal with inputs of random sizes, FusionVLAD [7] and AEGIS-Net [4] adopted more advanced 3D CNN with FusionVLAD focusing on cross-viewpoint VPR and AEGIS-Net concentrating on aggregating multi-level features.
@@
On the other hand, despite the fact that NetVLAD [2] is one of the most popular global descriptors, other global descriptors have also been explored. To start with, various forms of pooling have been employed for global descriptor generation due to their simplicity and generalizability. For example, APANet [33] favors sum pooling, Gordo et al. [34] chooses max pooling R-MAC [35], MinkLoc3D [36] adopts generalized-mean pooling (GeM) [37]. More recently, with Transformer [10] prevailing in a wide range of tasks, it has also been used to create global descriptors in VPR. For example, $R^{2}$ormer [11] employs attention weights to guide the aggregations of local tokens.
Although tremendous advances have been witnessed in VPR research, most of the work tackles the task under the batch learning setting with the complete dataset available prior to the training. However, our paper investigates the VPR task under a sequential learning setting, which is addressed by a few works.
@@
More related to our work, BioSLAM [38] tackles the VPR task from a lifelong learning perspective and chooses the generative replay from a rehearsal-based approach with a dual memory zone to store the previous observations. Nevertheless, AirLoop [18] combines a rehearsal-based approach with a regularization-based approach to mitigate the catastrophic forgetting issue with a lightweight model. Following this idea, InCloud [19] and CCL [22] extend to point cloud input with InCould introducing an angular-based knowledge distillation to relax the function regularization and CCL favoring contrast learning with InfoNCE [39] loss over metric learning with triplet loss [40]. Although we also build our work on AirLoop, our proposed VIPeR significantly differs from the above-mentioned methods with the adaptive mining for metric learning with triplet loss, a delicately designed three-zone memory bank for rehearsal, and probabilistic knowledge distillation (PKD) for regularization.
@@
Figure 2: Overview of main components and data flow within our proposed VIPeR. To alleviate the catastrophic forgetting in lifelong learning, our proposed VIPeR accompanies the place recognition model with a brain-inspired memory bank for rehearsal, and adaptive mining, relational memory aware synapses (RMAS) and probabilistic knowledge distillation (PKD) for regularization.
@@
As with many previous works on VPR, we cast the task of visual incremental place recognition as a retrieval problem. In particular, we define a sequence of environments =$\mathcal{E}=\{E_{0},E_{1},...,E_{t},...\}$ with only one environment being available for training at a time. Furthermore, we define a set of images as observations for each environment $E_{t}=\{I_{0},I_{1},...,I_{i},...\}$, which is also acquired in a sequential manner. Then, the VPR task is solved by incrementally constructing a database of place-tagged global descriptors $\mathcal{G}=\{\bm{g}_{0},...,\bm{g}_{i},...\}$ with each one generated with $\bm{g}_{i}=f(I_{i})$. $f(\cdot)$ here is a trainable model that consists of two sub-models, one extracts local features from input images and the other one aggregates the local features into global descriptors. Whenever the query image $I_{q}$ becomes available, the same model is used to generate a global descriptor from the query image $\bm{d}_{q}=f(I_{q})$. Following this, by matching the query descriptor to the database descriptors, we can retrieve the top-$K$ closest neighbors of the query image as the recognized places. In the following, we present VIPeR to tackle visual incremental place recognition by combining metric learning with lifelong learning. An overview of VIPeR is shown in Fig. 2.
@@ III-A Metric Learning
The naive triplet loss [40] is defined as
$L_{triplet}=\max\left(s^{an}_{j}-s^{ap}_{i}+\delta,0\right),$ (1)
@@
where $s^{an}_{j}=sim(f(I^{anc}),f(I^{neg}_{j}))$ is the cosine similarity between the global descriptors generated from the anchor image and a negative image sample, $s^{ap}_{i}$ is the cosine similarity between the ones from the anchor and a positive image sample, $\delta$ is a hyper-parameter indicating the margin. When selecting the positive and negative image, the naive triplet loss opted for a random mining strategy with the sample indices $i$ and $j$ randomly selected. Despite that, the naive triplet loss exhibits better generalizability and is used in previous works like AirLoop [18] and InCloud [19]; it also makes the training process less effective, resulting in less discriminative global descriptors.
@@
The hard triplet loss [41], to the contrary, goes to another extreme and favors the hard mining strategy. When choosing the sample images, the hard triplet loss focuses on the most difficult sample with the sample indices selected as $i=\text{argmin}_{i}(s^{ap}_{i})$ and $j=\text{argmax}_{j}(s^{an}_{j}$). Such a strategy largely increases the discrimination of the global descriptor and has been widely used by PointNetVLAD [31], AEGIS-Net [4] for place recognition in a single large-scale environment. However, such a mining strategy inevitably introduces bias into the training, leading to performance degradation in lifelong learning settings. We later demonstrate this in Table III in the ablation study.
To best join the advantages of random mining and hard mining, we propose a novel adaptive mining strategy and adjust the difficulty level of the training samples based on the fluctuations of the loss. We denote the triplet loss under the adaptive mining strategy as $L_{ada-triplet}$. Specifically, we keep tracking the changes in the loss during the training process. In the case when the loss increases in the current training step with the increment larger than a hyper-parameter $T_{d}$, we say the chosen training sample is too difficult, and we incline to a simpler sample in the following training steps. Contrarily, if the loss decreases in the current step with a decrement bigger than another hyper-parameter $T_{e}$, we determine the current training sample is too easy, and we lean towards more difficult ones. Under such a mining design, we achieve a balance between the discrimination requirements for place recognition and the generalizability requirements for handling the sequential input.
III-B Brain-inspired Memory Bank

### v3

# VIPeR: Visual Incremental Place Recognition with Adaptive Mining and Continual Learning

@@ Abstract
Visual place recognition (VPR) is essential to many autonomous systems. Existing VPR methods demonstrate attractive performance at the cost of limited generalizability. When deployed in unseen environments, these methods exhibit significant performance drops. Targeting this issue, we present VIPeR, a novel approach for visual incremental place recognition with the ability to adapt to new environments while retaining the performance of previous ones. We first introduce an adaptive mining strategy that balances the performance within a single environment and the generalizability across multiple environments. Then, to prevent catastrophic forgetting in continual learning, we design a novel multi-stage memory bank for explicit rehearsal. Additionally, we propose a probabilistic knowledge distillation to explicitly safeguard the previously learned knowledge. We evaluate our proposed VIPeR on three large-scale datasets—Oxford Robotcar, Nordland, and TartanAir. For comparison, we first set a baseline performance with naive finetuning. Then, several more recent continual learning methods are compared. Our VIPeR achieves better performance in almost all aspects with the biggest improvement of $13.85\%$ in average performance.
@@ I Introduction
Visual place recognition (VPR) aims to recognize a previously visited place by giving visual observations like images. It is a core component in simultaneous localization and mapping (SLAM) systems [1] to ensure the robust operation of autonomous agents in large-scale environments. When solving the VPR task, most of the existing approaches [2, 3, 4] cast it as a retrieval task. The solution then typically involves a two-step process, with step one generating global descriptors from local features and step two matching the descriptors to the ones in a place-tagged database.
Traditionally, VPR methods use hand-craft local features or semantic features and construct global descriptors with vector of locally aggregated descriptors (VLAD) [5], graph-based descriptors [6], etc. More recently, following the success of NetVLAD [2], there have been notable advancements in end-to-end VPR methods in terms of accuracy [3, 4] and robustness [7, 8]. However, these trained models have a strong independent and identical distribution (i.i.d.) assumption on the test data, the performance drops considerably when the test data is collected from an unseen environment that doesn’t follow the i.i.d. assumption.
@@
However, when deploying the VPR-equipped autonomous agents and AR/VR systems in real-world scenarios, it is difficult to collect sufficient data that covers all potential environments beforehand to pre-train the deep learning-based models. Such a gap between the advance in deep learning-based VPR methods and the practical needs of autonomous agents and AR/VR systems leads to the fact that recent practical-oriented systems still prefer the traditional VPR methods [9, 10]. In this paper, we aim to close this gap by enabling the continual learning ability of the deep learning-based VPR method. In particular, we present VIPeR, a visual incremental place recognition method that can adapt to the newly observed environment during deployment while reserving its knowledge of previously visited environments. Fig. 1 illustrates the core idea of the proposed VIPeR.
@@
The idea of enabling continual learning for deep neural networks is not new. Also known as continual learning or incremental learning, it is designed to overcome the notorious catastrophic forgetting when learning from sequential inputs. Despite that continual learning has been widely studied [11] and applied to many computer vision [12] or natural language processing [13] tasks, there is only a handful of works attempting to apply it to the place recognition task. AirLoop [14] and InCloud [15], as the pioneers in continual place recognition, both tackle the catastrophic forgetting with rehearsal-based and regularization-based methods.
@@
Although achieving promising performance, there is still plenty of room for improvement. As pointed out in CCL [16], the naive triplet loss used in AirLoop [14] and InCloud [15] inevitably introduces bias when learning the global representation of various places. Such a phenomenon becomes even severe when hard-mining is applied to the training. In addition, the naive choice of a simple memory structure substantially hinders the power of the rehearsal strategy. Finally, the deterministic knowledge distillation used in AirLoop and InCloud is very sensitive to the outliers in the data, leading to potential performance degradation.
@@
Targeting these issues, we first introduce an adaptive mining scheme for triplet loss in metric learning. Our adaptive mining scheme can adaptively choose hard mining or soft mining, achieving a balanced performance between a single environment and multiple environments. Focusing on continual learning, our memory bank differs from  [17, 18], achieves excellent results without any additional computational burden. Our memory bank is made of three components. For the current environment, we have a sensory memory bank and a working memory bank. For previously visited environments, we have another long-term memory bank. Finally, to make the model more robust against outliers, we extend the deterministic knowledge distillation used in AirLoop [14] and InCloud [15] to a probabilistic one. We conduct extensive experiments on three large-scale datasets and our proposed VIPeR demonstrates superior performance when compared to other state-of-the-art (SoTA) continual place recognition methods.
@@
  • We design a novel, multi-stage memory bank for rehearsal, alleviating the effects of catastrophic forgetting;
@@
II-A Continual Learning
Continual learning targets the sequential learning setting and focuses on improving the adaptability and generalizability of the deep learning models. However, naive solutions like finetuning on the new tasks make the model quickly forget the knowledge about previous tasks, resulting in catastrophic forgetting [19]. Addressing this issue, most existing methods can be categorized into rehearsal-based approaches and regularization-based approaches [12].
The core of the rehearsal-based approach is a memory buffer that either explicitly or implicitly stores samples from previous tasks. Then, when the model is trained on a new task, the samples stored in the buffer join the samples of the current task to refresh the model’s memory about previous tasks. Exemplary works in this approach are iCaRL [20], gradient episodic memory (GEM) [21], and deep generative replay (DGR) [22], which tackle the image classification task with iCaRL and GEM explicitly stores a small number of samples per class for rehearsal and DGR accompanies an additional generative model to enable implicit rehearsal. Building upon GEM, A-GEM [17] reduces the computational and memory costs by averaging the gradient of a random subset of episodic memory examples. Our memory bank, on the other hand, favors explicit memory storage and avoids introducing any additional computational overhead.
@@ II-B Learning-based Place Recognition
NetVLAD [2], a milestone in learning-based VPR methods, introduced differentiable VLAD [5] descriptors, enabling end-to-end training for VPR tasks. Building on this, SPE-VLAD [27] incorporated a spatial pyramid structure to encode structural information within the VLAD descriptor. Extending these advancements to 3D inputs, various methods employ 3D backbones, such as FusionVLAD [7], which focuses on cross-viewpoint VPR, and AEGIS-Net [4], which emphasizes aggregating multi-level features.
@@
In addition to NetVLAD [2], other pooling-based methods have also been explored for generating global descriptors, owing to their simplicity and generalizability. Examples include sum pooling in APANet [28], max pooling by Gordo et al. [29], and generalized-mean pooling (GeM) [30] adopted in MinkLoc3D [31]. More recently, with Transformer [32] prevailing in a wide range of tasks, it has also been used to create global descriptors in VPR, like $R^{2}$ormer [33].
To broaden the scope of this review, we acknowledge works like CosPlace [34] and EigenPlaces [35], which approach VPR as a classification problem. However, this paper focuses on VPR formulated as a retrieval problem.
With the advancement of visual foundation models, methods based on these pre-trained models have demonstrated compelling performance. Among them, CricaVPR [36] and SelaVPR [37] utilize pre-trained DINOv2 [38] and employ parameter-efficient transfer learning to adapt the model for VPR tasks. Focusing on global aggregation, SALAD [39] proposes a soft-assignment of DINOv2 features to NetVLAD’s clusters while SegVLAD [40] augments NetVLAD descriptor with supersegments from SAM [41]. Finally, AnyLoc [42] and VDNA-PR [43] both explore the data representation power of the visual foundation model. In particular, VDNA-PR combines the features extracted by DINOv2 with the visual distribution of neuron activations representation, but AnyLoc experiments with a wide range of data, from indoor to outdoor, aerial to underwater. Although the use of visual foundation models reduces the likelihood of encountering out-of-distribution data, continual learning is still indispensable, as we demonstrate in Table II.
@@
Despite tremendous advances in VPR research, most of the work tackles the task under the batch learning setting with the complete dataset available prior to the training. However, our paper investigates the VPR task under a sequential learning setting, which is addressed by a few works. More related to our work, BioSLAM [18] tackles the VPR task from a continual learning perspective and chooses the generative replay from a rehearsal-based approach with a dual memory zone to store the previous observations. Nevertheless, AirLoop [14] combines a rehearsal-based approach with a regularization-based approach to mitigate the catastrophic forgetting issue with a lightweight model. Following this idea, InCloud [15] and CCL [16] extend to point cloud input with InCould introducing an angular-based knowledge distillation to relax the function regularization and CCL favoring contrast learning with InfoNCE [44] loss over metric learning with triplet loss [45]. Our VIPeR significantly differs from the above-mentioned methods with the adaptive mining for metric learning with triplet loss, a delicately designed multi-stage memory bank for rehearsal, and probabilistic knowledge distillation (PKD) for regularization.
@@
Figure 2: Overview of main components and data flow within our proposed VIPeR. To alleviate the catastrophic forgetting in continual learning, our proposed VIPeR accompanies the place recognition model with a multi-stage memory bank for rehearsal, and adaptive mining, relational memory aware synapses (RMAS) and probabilistic knowledge distillation (PKD) for regularization.
@@
As with many previous works on VPR, we cast the task of visual incremental place recognition as a retrieval problem. In particular, we define a sequence of environments =$\mathcal{E}=\{E_{0},E_{1},...,E_{t},...\}$ with only one environment being available for training at a time. Furthermore, we define a set of images as observations for each environment $E_{t}=\{I_{0},I_{1},...,I_{i},...\}$, which is also acquired in a sequential manner. Then, the VPR task is solved by incrementally constructing a database of place-tagged global descriptors $\mathcal{G}=\{\bm{g}_{0},...,\bm{g}_{i},...\}$ with each one generated with $\bm{g}_{i}=f(I_{i})$. $f(\cdot)$ here is a trainable model that consists of two sub-models, one extracts local features from input images and the other one aggregates the local features into global descriptors. Whenever the query image $I_{q}$ becomes available, the same model is used to generate a global descriptor from the query image $\bm{d}_{q}=f(I_{q})$. Following this, by matching the query descriptor to the database descriptors, we can retrieve the top-$K$ closest neighbors of the query image as the recognized places. In the following, we present VIPeR to tackle visual incremental place recognition by combining metric learning with continual learning. An overview of VIPeR is shown in Fig. 2.
@@ III-A Metric Learning
The naive triplet loss [45] is defined as $L_{triplet}=\max\left(s^{an}_{j}-s^{ap}_{i}+\delta,0\right)$, where $s^{an}_{j}=sim(f(I^{anc}),f(I^{neg}_{j}))$ is the cosine similarity between the global descriptors generated from the anchor image and a negative image sample, $s^{ap}_{i}$ is the cosine similarity between the ones from the anchor and a positive image sample, $\delta$ is a hyper-parameter indicating the margin. When selecting the positive and negative image, the naive triplet loss opted for a random mining strategy with the sample indices $i$ and $j$ randomly selected. Despite that, the naive triplet loss exhibits better generalizability and is used in previous works like AirLoop [14] and InCloud [15]; it also makes the training process less effective, resulting in less discriminative global descriptors.
@@
The hard triplet loss [46], to the contrary, goes to another extreme and favors the hard mining strategy. When choosing the sample images, the hard triplet loss focuses on the most difficult sample with the sample indices selected as $i=\text{argmin}_{i}(s^{ap}_{i})$ and $j=\text{argmax}_{j}(s^{an}_{j}$). Such a strategy largely increases the discrimination of the global descriptor and has been widely used [4, 47] for place recognition in a single large-scale environment. However, such a mining strategy inevitably introduces bias into the training, leading to performance degradation in continual learning settings. We later demonstrate this in Table III.
To balance between the hard and random mining, we introduce adaptive mining. Unlike existing adaptive mining methods [48, 49] that only adjust negative samples, our method dynamically choose different positive and negative samples. Especially, we use the fluctuation of the loss as the criterion to determine the difficulty:
$\Delta L_{triplet}=\lVert s_{j}^{an,k}-s_{i}^{ap,k}-s_{j}^{an,k-1}+s_{i}^{ap,k+1}\rVert_{1}$ (1)
where $k$ denotes the data at the $k$-th step. When the loss increment is greater than $T_{d}$, we consider the current training difficulty to be too high, and continued training may lead to insufficient generalization performance of the model. Therefore, with $s_{j}^{an}$ and $s_{i}^{ap}$ sorted in descending and ascending orders, we decrease $i$ and increase $j$ to reduce the training difficulty. Contrarily, when the loss decrement is greater than $T_{e}$, we increase $i$ and decrease $j$ to increase the training difficulty. By adjusting the training difficulty through $\Delta L_{triplet}$, both positive and negative samples can be considered, ensuring the model’s generalization ability.
III-B Multi-stage Memory Bank